OpenStudy (anonymous):

Show that the equation represents a circle by rewriting it in standard form. 4x2 + 4y2 + 8x − y = 0

OpenStudy (whpalmer4):

Complete the square on x and y respectively. Might want to divide through by 4 to eliminate the coefficients of x^2 and y^2 first.

OpenStudy (anonymous):

divide it first y 4 so that the coefficient of x^2 and y^2 will be one

OpenStudy (anonymous):

but i won't have a radius since it =0, so then what?

OpenStudy (whpalmer4):

Gotta have faith :-)

OpenStudy (whpalmer4):

You will by the time the dust settles on the algebra.

OpenStudy (anonymous):

faith isn't going to finish my hw-_- haha

OpenStudy (whpalmer4):

Seriously, when you complete the square, don't you end up adding stuff to both sides? That's going to give you your radius.

OpenStudy (anonymous):

oooooh i see now!

OpenStudy (whpalmer4):

Got an answer for me to check?

OpenStudy (anonymous):

(x-1)^2+(y+1/8)^2=65/64 ?

OpenStudy (whpalmer4):

Very close! \[(x+1)^2 + (y-\frac{1}8)^2 = \frac{65}{64}\]

OpenStudy (whpalmer4):

You should look at your work again and figure out how you got your signs switched around. Here's what I did: \[4x^2+4y^2+8x-y=0\]Divide through by 4\[x^2+y^2+2x-\frac{y}{4} = 0\]Rearrange the \(x\) stuff first: \[(x^2+2x) + (y^2 - \frac{y}4) = 0\]We take half of the coefficient of \(x\), square it, and add to both sides: \[(x^2+2x+(\frac{2}{2})^2) + (y^2-\frac{y}4) = (\frac{2}{2})^2\]Rewrite the \(x\) terms as a perfect square\[(x+1)^2 + (y^2-\frac{y}4) = 1\]Take half of the coefficient of \(y\), square it, and add to both sides:\[(x+1)^2 + (y^2-\frac{y}4 + (-\frac{1}{8})^2) = 1+(-\frac{1}8)^2\]Rewrite the \(y\) terms as a perfect square:\[(x+1)^2 + (y-\frac{1}{8})^2 = 1+\frac{1}{64}\]

OpenStudy (anonymous):

oh okay i found where i mixed them. thank you so much! can you help me with another? how would I find the area of this triangle if it isn't a right triangle?

OpenStudy (whpalmer4):

Area of a triangle is given by \[A = \frac{1}{2}bh\]where \(b\) is the length of the base and \(h\) is the height. Works even for a weird triangle like this one. What's the length of the base? What's the height?

OpenStudy (anonymous):

so 1/2*8*3?

OpenStudy (whpalmer4):

Yep. And you could also do it by "extending" the triangle so that it had a vertex at (7,1) instead of at B, find the area of that triangle (1/2 * 9*3 = 27/2 = 13.5), then subtract off the little triangle added by doing so (1/2 * 1*3 = 1.5) and you get the same answer.

OpenStudy (anonymous):

oh okay! okay so for this one, the equation is x2/9+ y2/4= 1

OpenStudy (anonymous):

Find the x- and y-intercepts.

OpenStudy (whpalmer4):

Yes, that's the equation for that ellipse. The x and y intercepts are found by setting y and x = 0 respectively and solving for the remaining variable.

OpenStudy (whpalmer4):

My cat is demanding that I go to bed now, it's cold in the house and she wants a warm body to snuggle with :-)

OpenStudy (anonymous):

nooo don't leave me! can you help me with one more pweeeease

OpenStudy (whpalmer4):

Okay, it'll have to be quick :-)

OpenStudy (anonymous):

okay! simplify the compound fraction: 9x-9y/x/y+y/x

OpenStudy (whpalmer4):

Can you write that with parentheses so I can tell what it is supposed to be?

OpenStudy (anonymous):